# New Integrable Chiral Cosmological Models with Two Scalar Fields

### v1

@@ I Introduction
Most cosmological models, which describe the global evolution of the Universe, include scalar fields. To describe only one epoch of the Universe evolution, single field models can be used. Using two-field models, one can describe multiple epochs or multiple physical phenomenons. Models with a single scalar field nonminimally coupled to gravity can be transformed to models with a minimally coupled scalar field with a canonical kinetic term by the metric and scalar field transformations. On the other hand, it is not possible to transform a model with two scalar fields nonminimally coupled to gravity to a model with two minimally coupled scalar fields and a standard kinetic part of the Lagrangian in the most general case [1]. After the metric transformation, one obtains a general relativity model with nonstandard kinetic terms of scalar fields, a so-called chiral cosmological model (CCM) [2].
@@
In the Friedmann–Lemaître–Robertson–Walker (FLRW) metric, the Einstein’s equations reduce to a system of ordinary differential equations. The problem of searching for integrable cosmological models with scalar fields is being actively investigated, essentially in the case of single-field cosmological models, but such models have been found only for a few specific forms of the scalar field potential [3, 4, 5, 6, 7, 8, 9, 10, 11].
@@
A two-field CCM with a constant potential is trivially integrated not only in the spatially-flat FLRW metric, but also in the Bianchi I metric [12, 13]. At the same time, finding integrable cosmological models with multiple scalar fields is, in general, an extremely complicated task. We know only a few examples of integrable cosmological models with two or more fields [14, 15, 16, 17, 18]. The goal of this paper is to construct new integrable CCM with two fields and potentials written in terms of hyperbolic functions.
@@
There are a few methods to integrate evolution equations in the FLRW metric. The integrability can be proven by using a suitable parametric time [6], but it is not immediately clear as to how to find this parametric time. The superpotential method, which is also known as the Hamilton-Jacobi equation approach, is useful to integrate not only single-field models with the exponential potential [3], but also two-field models with the exponential potential [18]. Note that this method is actively used to study inflation and to construct cosmological models with particular exact solutions [19, 3, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 18, 35].
@@
Another method for constructing one-field integrable models is based on the conformal metric transformation and the construction of the corresponding model in the Jordan frame [5, 8, 10, 36]. In Ref. [5], the authors construct two-field models with one ordinary scalar field, one phantom field and induced gravity terms to prove the integrability of a single-field cosmological model with a potential written in terms of hyperbolic functions. Note that integrability of this model has been proven in the closed and open Friedmann universes as well [37]. In Refs. [5, 38, 39], the proposed cosmological model has been intensively studied in the context of the early Universe evolution, with special emphasis on inflation and the possibility of crossing the Big Bang-Big Crunch singularity. In Ref. [10], the authors have proven integrability of the same model in a simpler way by using a single-field integrable model with non-minimal coupling constructed in Ref. [9]. Note that this model belongs to a wide class of integrable models with minimally coupled scalar fields and potentials in terms of hyperbolic functions. Their integrability have been proven by using a suitable parametric time in Ref. [6] (see also Ref. [10]).
@@
To construct a new integrable CCM with two fields, we start from models with two nonminimally coupled scalar fields. The single-field integrable model proposed in Ref. [9] has an interesting feature: the Ricci scalar is an integral of motion. Recently, $N$-field cosmological models with the same property have been found and their integrability in the spatially flat FLRW metric have been proven [40]. In this paper, we find general solutions of evolution equations in the Friedmann universe with arbitrary spatial curvature for such two-field integrable models with non-minimal coupling. After finding such solutions, we obtain two-field chiral cosmological models in the Einstein frame by the conformal transformation of the metric. For a few two-field potentials, we get general solutions. In the conformal time, we get analytic expressions of the general solutions in terms of the Jacobi elliptic functions.
@@ II Two-field model with nonminimal coupling
$T_{\mu\nu}=C_{1}\partial_{\mu}\sigma^{1}\partial_{\nu}\sigma^{1}+C_{2}\partial_{\mu}\sigma^{2}\partial_{\nu}\sigma^{2}-g_{\mu\nu}\left(\frac{1}{2}g^{\alpha\beta}\left(C_{1}\partial_{\alpha}\sigma^{1}\partial_{\beta}\sigma^{1}+C_{2}\partial_{\alpha}\sigma^{2}\partial_{\beta}\sigma^{2}\right)+V\right)\,.$
@@
$A=1,2$; $F_{,\sigma^{A}}\equiv\frac{dF}{d\sigma^{A}}$ for any function $F(\sigma^{1},\sigma^{2})$, and there is no summation in $A$.
From Eq. (2), it follows
@@
With such a function $U$, Eq. (6) is
$2V+\sum\limits_{A=1}^{2}\frac{1}{2}\sigma^{A}V_{,\sigma^{A}}=RU_{0}.$ (10)
@@ III The corresponding models in the Einstein frame
Now we construct models in the Einstein frame that correspond to the previously chosen functions $U$ and $V$. To get a model with two standard scalar fields, we take $C_{A}=1$. Also, we can take $\sigma^{A}_{0}=0$ without loss of generality.
@@
$g^{\mathrm{E}}_{\mu\nu}=\frac{2U}{M^{2}_{\mathrm{Pl}}}g_{\mu\nu},$ (12)
@@
$S_{\mathrm{E}}=\int d^{4}x\sqrt{-g^{\mathrm{E}}}\left(\frac{M^{2}_{\mathrm{Pl}}}{2}R_{\mathrm{E}}-\frac{g_{\mathrm{E}}^{\mu\nu}}{2}\sum\limits_{A,B=1}^{2}K_{AB}\partial_{\mu}\sigma^{A}\partial_{\nu}\sigma^{B}-V_{\mathrm{E}}\right),$ (13)
@@
By varying the action (20), we get the following equations:
@@
$\Box\chi^{1}-\frac{\sqrt{6}}{12M_{\mathrm{Pl}}}\sinh\left(\frac{\chi^{1}}{3\sqrt{6}M_{\mathrm{Pl}}}\right)g^{(\mathrm{E})\mu\nu}\partial_{\mu}\chi^{2}\partial_{\nu}\chi^{2}-\frac{\partial V_{\mathrm{E}}}{\partial\chi^{1}}=0;$ (23)
$\cosh^{2}\left(\frac{\chi^{1}}{\sqrt{6}M_{\mathrm{Pl}}}\right)\Box\chi^{2}+\frac{\sqrt{6}}{6M_{\mathrm{Pl}}}\sinh\left(\frac{\chi^{1}}{3\sqrt{6}M_{\mathrm{Pl}}}\right)g^{(\mathrm{E})\mu\nu}\partial_{\mu}\chi^{1}\partial_{\nu}\chi^{2}-\frac{\partial V_{\mathrm{E}}}{\partial\chi^{2}}=0,$ (24)
@@
-- $g^{(\mathrm{E})}_{\mu\nu}\left(\frac{1}{2}g^{\mathrm{(E)\alpha\beta}}\partial_{\alpha}\chi^{1}\partial_{\beta}\chi^{2}+\frac{1}{2}\cosh^{2}\left(\frac{\chi^{1}}{\sqrt{6}M_{\mathrm{Pl}}}\right)g^{(\mathrm{E})\alpha\beta}\partial_{\alpha}\chi^{2}\partial_{\beta}\chi^{2}+V_{\mathrm{E}}(\chi^{1},\chi^{2})\right).$
In the Einstein frame, the d’Alembert operator $\Box$ acting on a scalar is
$\Box=\frac{1}{\sqrt{g^{(\mathrm{E})}}}\partial_{\mu}\left(\sqrt{g^{(\mathrm{E})}}g^{(\mathrm{E})\mu\nu}\partial_{\nu}\right).$ (26)
@@ IV.1 The FLRW metric with conformal time
In this metric, Eqs. (2) and (3) with $C_{A}=1$ have the following form:
@@
$\ddot{\sigma}^{A}+2h\dot{\sigma}^{A}-6U_{,\sigma^{A}}^{\prime}a^{2}R+a^{2}V_{,\sigma^{A}}=0,$ (30)
where $h=\dot{a}/a$, and “dots” denote derivatives with respect to the parametric time $\tau$.
@@
$a(\tau)=\left\{\begin{aligned} a_{1}\sin(\sqrt{K}\tau)+a_{2}\cos(\sqrt{K}\tau),\quad&\quad K>0;\\ a_{1}\tau+a_{0},\qquad\qquad\qquad\quad\quad\quad\,\,\,&\quad K=0;\\ a_{1}\sinh(\sqrt{K}\tau)+a_{2}\cosh(\sqrt{K}\tau),&\quad K<0;\\ \end{aligned}\right.$ (35)
where $a_{0}$, $a_{1}$ and $a_{2}$ are arbitrary constants.
@@
System (37) has the first integral
@@
Using Eq. (38), we get from Eq. (40) the following equation:
@@
$t-t_{0}=\pm{\frac{\sqrt{3}}{\sqrt{R_{0}}}\ln\left({\frac{R_{0}\,U_{0}{a}^{2}-6U_{0}K}{\sqrt{R_{0}}}}+\sqrt{R_{0}\,U_{0}^{2}{a}^{4}-12KU_{0}^{2}{a}^{2}+2EU_{0}\,}\right)}.$ (42)
At $R_{0}=0$, we obtain the following nonconstant solutions:
$a(t)=\left\{\begin{split}&\pm\frac{\sqrt{K(E-6K^{2}U_{0}^{3}(t-t_{0})^{2})}}{\sqrt{6}U_{0}K},\quad K\neq 0,\\ &\pm\frac{\sqrt{3}}{3U_{0}}\sqrt{\pm\sqrt{6EU_{0}\,}U_{0}(t-t_{0})\,},\quad K=0\,.\end{split}\right.$ (43)
Equation (41) coincides with the equation obtained in Ref. [9] for $K=0$ and in Ref. [10] for an arbitrary $K$ in the case of the corresponding single-field model. Thus, the analysis of the Universe evolution in depending on the values of the constants $K$, $R_{0}$ and $E$, including conditions of the existence of bounce solutions given in Ref. [10], is also valid for the considered two-field models. For this reason, we do not repeat it in this paper. Note only that in the case $K=0$, the Hubble parameter has the following simple form:
$H(t)=\left\{\begin{split}&\frac{\sqrt{3R_{0}}\left(\mathrm{e}^{2\sqrt{3R_{0}}t/3}+B\right)}{6\left(\mathrm{e}^{2\,\sqrt{3R_{0}}t/3}-B\right)}\,,\quad R_{0}>0\\ &\frac{1}{2(t-t_{0})},\quad R_{0}=0\,.\end{split}\right.$ (44)
@@
Note that explicit form of the $a(\tau)$ does not depend on the form of the potential and is given by Eq. (34) for $R_{0}>0$ and Eq. (35) for $R_{0}=0$. To get the general solutions in the conformal time, so it is suitable to write out expressions for $H(\tau)$ and $a(\tau)$ in explicitly real forms. We present formulas for the case of $K=0$ and $R_{0}>0$ only.
For $K=0$ and $R_{0}>0$, we obtain
@@
    $a_{0}\sqrt{\frac{R_{0}}{12}}\left(\tau-\tau_{0}\right)\in\left(0,K\left(\frac{1}{2}\right)\right);$
@@
    $a_{0}\sqrt{\frac{R_{0}}{6}}\left(\tau-\tau_{0}\right)\in\left(-K\left(\frac{1}{2}\right),K\left(\frac{1}{2}\right)\right);$
@@
$K(m)\equiv F(\pi/2\left|m\right.)$ is the complete elliptic integral of the first kind, and the Jacobi elliptic functions $\mathrm{sn}$, $\mathrm{cn}$ and $\mathrm{dn}$ have been used.

### v3

@@ I Introduction
Many cosmological models, which describe the global evolution of the Universe, include scalar fields. To describe only one epoch of the Universe evolution, single-field models can be used. Using two-field models, one can explore multiple epochs or multiple physical phenomena. Models with a single scalar field nonminimally coupled to gravity can be transformed to models with a minimally coupled scalar field with a canonical kinetic term by metric and scalar field transformations. On the other hand, it is not possible to transform a model with two scalar fields nonminimally coupled to gravity to a model with two minimally coupled scalar fields and a standard kinetic part of the Lagrangian in the most general case [1]. After the metric transformation, one obtains a general relativity model with nonstandard kinetic terms of scalar fields, a so-called chiral cosmological model (CCM) [2].
@@
In the Friedmann-Lemaître-Robertson-Walker (FLRW) metric, the Einstein’s equations reduce to a system of ordinary differential equations. The problem of searching for integrable cosmological models with scalar fields is being actively investigated, essentially in the case of single-field cosmological models, but such models have been found only for a few specific forms of the scalar field potential [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
@@
A two-field CCM with a constant potential is trivially integrated not only in the spatially flat FLRW metric, but also in the Bianchi I metric [20, 21]. At the same time, the search for integrable cosmological models with multiple scalar fields and nonconstant potentials is an extremely complicated task. We know only a few examples of integrable cosmological models with two or more fields [22, 23, 17, 24, 25, 26]. The goal of this paper is to construct new integrable CCMs with two scalar fields and potentials written in terms of hyperbolic functions.
@@
There are a few methods to integrate evolution equations in the FLRW metric. The integrability can be proven by using a suitable parametric time [10], but it is not immediately clear as to how to find this parametric time. The superpotential method, which is also known as the Hamilton-Jacobi equation approach, is useful to integrate not only single-field models with the exponential potential [3], but also two-field models with the exponential potential [26]. Note that this method is actively used to study inflation and to construct cosmological models with particular exact solutions [27, 3, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 26, 44].
@@
Another method for constructing single-field integrable models is based on the conformal metric transformation and the construction of the corresponding model in the Jordan frame [7, 12, 14, 45]. In Ref. [7], the authors construct two-field models with one ordinary scalar field, one phantom field, and induced gravity terms to prove the integrability of a single-field cosmological model with a potential written in terms of hyperbolic functions. Note that integrability of this model has been proven in the closed and open Friedmann universes as well [9]. In Refs. [7, 46, 47], the proposed cosmological model has been intensively studied in the context of the early Universe evolution, with special emphasis on inflation and the possibility of crossing the big bang-big crunch singularity. In Ref. [14], the authors have proven integrability of the same model in a simpler way by using a single-field integrable model with nonminimal coupling constructed in Ref. [13]. Note that this model belongs to a wide class of integrable models with minimally coupled scalar fields and potentials in terms of hyperbolic functions. Their integrability have been proven by using a suitable parametric time in Ref. [10] (see also Ref. [14]).
@@
To construct a new integrable CCM with two fields, we start from models with two nonminimally coupled scalar fields. The single-field integrable model proposed in Ref. [13] has an interesting feature: the Ricci scalar is an integral of motion. Recently, $N$-field cosmological models with the same property have been found and their integrability in the spatially flat FLRW metric have been proven [48].
In this paper, we find general solutions of evolution equations in the Friedmann universe with arbitrary spatial curvature for a few such two-field integrable models with nonminimal coupling. We show that in the conformal time, the system of the evolution equations can be transformed to a Hamiltonian system of two equations. For a few two-field polynomial potentials, we integrate this system and get analytic expressions of the general solutions in terms of the Jacobi elliptic functions. After finding such solutions, we obtain two-field chiral cosmological models in the Einstein frame by the conformal transformation of the metric. The obtained integrable CCMs have potentials represented in terms of hyperbolic functions.
@@ II Two-field model with nonminimal coupling
$T_{\mu\nu}=C_{1}\partial_{\mu}\sigma^{1}\partial_{\nu}\sigma^{1}+C_{2}\partial_{\mu}\sigma^{2}\partial_{\nu}\sigma^{2}-g_{\mu\nu}\left(\frac{1}{2}g^{\alpha\beta}\left(C_{1}\partial_{\alpha}\sigma^{1}\partial_{\beta}\sigma^{1}+C_{2}\partial_{\alpha}\sigma^{2}\partial_{\beta}\sigma^{2}\right)+V\right)\,,$
@@
$A=1,2$; $\,F_{,\sigma^{A}}\equiv\frac{\partial F}{\partial\sigma^{A}}$ for any function $F(\sigma^{1},\sigma^{2})$, and there is no summation in $A$.
From Eq. (2), it follows that
@@
With such a function $U$, Eq. (6) takes the form
$2V-\frac{1}{2}\sum\limits_{A=1}^{2}\sigma^{A}V_{,\sigma^{A}}=RU_{0}.$ (10)
@@ III The corresponding models in the Einstein frame
Now we construct models in the Einstein frame that correspond to the previously chosen functions $U$ and $V$. To get a model with two standard scalar fields, we take all $C_{A}=1$. Also, we can take $\sigma^{A}_{0}=0$ without loss of generality.
@@
$g^{(\mathrm{E})}_{\mu\nu}=\frac{2U}{M^{2}_{\mathrm{Pl}}}g_{\mu\nu},$ (12)
@@
$S_{\mathrm{E}}=\int d^{4}x\sqrt{-g^{(\mathrm{E})}}\left(\frac{M^{2}_{\mathrm{Pl}}}{2}R_{\mathrm{E}}-\frac{g^{(\mathrm{E})\mu\nu}}{2}\sum\limits_{A,B=1}^{2}K_{AB}\partial_{\mu}\sigma^{A}\partial_{\nu}\sigma^{B}-V_{\mathrm{E}}\right),$ (13)
@@
By varying action (20), we get the following equations:
@@
$\Box^{(\mathrm{E})}\chi^{1}-\frac{\sqrt{6}}{12M_{\mathrm{Pl}}}\sinh\left(\frac{\chi^{1}}{3\sqrt{6}M_{\mathrm{Pl}}}\right)g^{(\mathrm{E})\mu\nu}\partial_{\mu}\chi^{2}\partial_{\nu}\chi^{2}-\frac{\partial V_{\mathrm{E}}}{\partial\chi^{1}}=0;$ (23)
$\cosh^{2}\left(\frac{\chi^{1}}{\sqrt{6}M_{\mathrm{Pl}}}\right)\Box^{(\mathrm{E})}\chi^{2}+\frac{\sqrt{6}}{6M_{\mathrm{Pl}}}\sinh\left(\frac{\chi^{1}}{3\sqrt{6}M_{\mathrm{Pl}}}\right)g^{(\mathrm{E})\mu\nu}\partial_{\mu}\chi^{1}\partial_{\nu}\chi^{2}-\frac{\partial V_{\mathrm{E}}}{\partial\chi^{2}}=0,$ (24)
@@
-- $g^{(\mathrm{E})}_{\mu\nu}\left(\frac{1}{2}g^{\mathrm{(E)\alpha\beta}}\partial_{\alpha}\chi^{1}\partial_{\beta}\chi^{1}+\frac{1}{2}\cosh^{2}\left(\frac{\chi^{1}}{\sqrt{6}M_{\mathrm{Pl}}}\right)g^{(\mathrm{E})\alpha\beta}\partial_{\alpha}\chi^{2}\partial_{\beta}\chi^{2}+V_{\mathrm{E}}(\chi^{1},\chi^{2})\right).$
The d’Alembert operator $\Box^{(\mathrm{E})}$ acting on a scalar is
$\Box^{(\mathrm{E})}=\frac{1}{\sqrt{-g^{(\mathrm{E})}}}\partial_{\mu}\left(\sqrt{-g^{(\mathrm{E})}}g^{(\mathrm{E})\mu\nu}\partial_{\nu}\right).$ (26)
@@ IV.1 The FLRW metric with conformal time
In this metric, Eqs. (2) and (3) with $C_{A}=1$ have the following forms:
@@
$\ddot{\sigma}^{A}+2h\dot{\sigma}^{A}-6U_{,\sigma^{A}}a^{2}R+a^{2}V_{,\sigma^{A}}=0,$ (30)
where $h=\dot{a}/a$, and “dots” denote derivatives with respect to the conformal time $\tau$.
@@
$a(\tau)=\left\{\begin{aligned} a_{1}\sin(\sqrt{K}\tau)+a_{2}\cos(\sqrt{K}\tau),\quad&\quad K>0;\\ a_{1}\tau+a_{0},\qquad\qquad\qquad\quad\quad\quad\,\,\,&\quad K=0;\\ a_{1}\sinh(\sqrt{-K}\tau)+a_{2}\cosh(\sqrt{-K}\tau),&\quad K<0;\\ \end{aligned}\right.$ (35)
where $a_{0}$, $a_{1}$, and $a_{2}$ are arbitrary constants.
@@
System (37) is a Hamiltonian one with
${\cal{H}}=\frac{1}{2}\left(p^{1}\right)^{2}+\frac{1}{2}\left(p^{2}\right)^{2}+\frac{K}{2}\left[\left(y^{1}\right)^{2}+\left(y^{2}\right)^{2}\right]+\tilde{V},$ (38)
where $p_{i}$ are momenta. It has the first integral
@@
Using Eq. (39), we transform Eq. (41) to the following equation:
@@
$a^{2}(t)=\frac{36K^{2}U_{0}^{2}-2EU_{0}R_{0}+12KU_{0}\sqrt{R_{0}}\,{\rm e}^{\pm\sqrt{3R_{0}}\left(t-t_{0}\right)/3}+R_{0}\,{\rm e}^{\pm 2\sqrt{3R_{0}}\left(t-t_{0}\right)/3}}{2U_{0}R_{0}^{3/2}{\rm e}^{\pm\sqrt{3R_{0}}\left(t-t_{0}\right)/3}}\,,$ (43)
where $t_{0}$ is a constant. So, the Hubble parameter is
$H(t)=\pm\,\frac{\sqrt{3R_{0}}\left(2EU_{0}R_{0}-36K^{2}U_{0}^{2}+R_{0}\,{\rm e}^{\pm\sqrt{3R_{0}}\left(t-t_{0}\right)/3}\right)}{6\left(36K^{2}U_{0}^{2}+12KU_{0}\sqrt{R_{0}}\,{\rm e}^{\pm\sqrt{3R_{0}}\left(t-t_{0}\right)/3}+R_{0}\,{\rm e}^{\pm 2\sqrt{3R_{0}}\left(t-t_{0}\right)/3}-2U_{0}R_{0}E\right)}\,.$ (44)
In the case of $K=0$, the Hubble parameter has the following simple form:
$H(t)=\frac{\sqrt{3R_{0}}\left(\mathrm{e}^{2\sqrt{3R_{0}}t/3}+B\right)}{6\left(\mathrm{e}^{2\,\sqrt{3R_{0}}t/3}-B\right)}\,,$ (45)
@@
For $R_{0}=0$, one gets
$a^{2}(t)=\left\{\begin{split}&\frac{E-6K^{2}U_{0}(t-t_{0})^{2}}{6U_{0}K},\quad K\neq 0,\\ &\frac{\sqrt{6EU_{0}\,}}{3U_{0}}|t-t_{0}|\,,\quad K=0\,.\end{split}\right.$ (46)
The Hubble parameter is
$H(t)=\left\{\begin{split}&\frac{6K^{2}U_{0}\left(t-t_{0}\right)}{6K^{2}U_{0}\left(t-t_{0}\right)^{2}-E},\quad K\neq 0\\ &\frac{1}{2(t-t_{0})},\quad K=0\,.\end{split}\right.$ (47)
Equation (42) coincides with the equation obtained in Ref. [13] for $K=0$ and in Ref. [14] for an arbitrary $K$ in the case of the corresponding single-field model. The detail analysis of the Universe evolution in depending on the values of the constants $K$, $R_{0}$, and $E$, with conditions of the existence of bounce solutions, is given in Ref. [14] without explicit expressions for $a(t)$ and $H(t)$. We do not repeat it in this paper, because it is also valid for the considered two-field models.
Note that the explicit form of the $a(\tau)$ does not depend on the form of the potential and is given by Eq. (34) for $R_{0}>0$ and Eq. (35) for $R_{0}=0$. To get the general solutions in the conformal time, it is suitable to use expressions for $H(\tau)$ and $a(\tau)$ in explicitly real forms.
We present formulas for the case of $K=0$ and $R_{0}>0$:
@@
    $a_{0}\sqrt{\frac{R_{0}}{12}}\left(\tau-\tau_{0}\right)\in\left(0,K\left(\frac{1}{2}\right)\right).$
@@
    $a_{0}\sqrt{\frac{R_{0}}{6}}\left(\tau-\tau_{0}\right)\in\left(-K\left(\frac{1}{2}\right),K\left(\frac{1}{2}\right)\right).$
@@
$K(m)\equiv F(\pi/2\left|m\right.)$ is the complete elliptic integral of the first kind, and the Jacobi elliptic functions $\mathrm{sn}$, $\mathrm{cn}$, and $\mathrm{dn}$ have been used.
Solving Eq. (42) with an arbitrary $K$, one obtains:
  1. 1.
    at $\alpha^{2}>0$, $E/U_{0}>0$, $K>0$:
    $a(\tau)=\sqrt{\alpha+\frac{6K}{R_{0}}}\mathrm{dc}\left(\left.\sqrt{\alpha\frac{R_{0}}{12}+2K}(\tau-\tau_{0})\right|\frac{6K-\alpha R_{0}}{6K+\alpha R_{0}}\right);$ (54)
  2. 2.
    at $\alpha^{2}>0$, $E/U_{0}>0$, $K<0$:
    $a(\tau)=\sqrt{-\alpha+\frac{6|K|}{R_{0}}}\mathrm{sc}\left(\left.\sqrt{\alpha\frac{R_{0}}{12}+2|K|}(\tau-\tau_{0})\right|\frac{2\alpha R_{0}}{6|K|+\alpha R_{0}}\right);$ (55)
  3. 3.
    at $\alpha^{2}>0$, $E/U_{0}<0$:
    $a(\tau)=\sqrt{\alpha+\frac{6K}{R_{0}}}\mathrm{nc}\left(\left.\sqrt{\alpha\frac{R_{0}}{6}}(\tau-\tau_{0})\right|\frac{1}{2}-\frac{3K}{R_{0}\alpha}\right);$ (56)
  4. 4.
    at $\alpha^{2}<0$:
    $a(\tau)=\left(\frac{2E}{R_{0}U_{0}}\right)^{1/4}\frac{\mathrm{sn}\left(\left.\sqrt{\frac{R_{0}}{3}}\left(\frac{2E}{R_{0}U_{0}}\right)^{1/4}(\tau-\tau_{0})\right|\frac{1}{2}+\frac{3K}{R_{0}}\sqrt{\frac{R_{0}U_{0}}{2E}}\right)}{1+\mathrm{cn}\left(\left.\sqrt{\frac{R_{0}}{3}}\left(\frac{2E}{R_{0}U_{0}}\right)^{1/4}(\tau-\tau_{0})\right|\frac{1}{2}+\frac{3K}{R_{0}}\sqrt{\frac{R_{0}U_{0}}{2E}}\right)}.$ (57)
Here
$\alpha=\sqrt{\left(\frac{6K}{R_{0}}\right)^{2}-\frac{2E}{R_{0}U_{0}}},\quad\mathrm{dc}=\frac{\mathrm{dn}}{\mathrm{cn}},\quad\mathrm{sc}=\frac{\mathrm{sn}}{\mathrm{cn}},\quad\mathrm{nc}=\frac{1}{\mathrm{cn}}.$
In the two last cases $K$ is arbitrary.
Also, for the special case $E=0$, $K=0$, one has
$a(\tau)=\frac{a_{0}}{1\pm a_{0}\sqrt{\frac{R_{0}}{12}}(\tau-\tau_{0})}.$ (58)
V Potentials and solutions of the field equations
V.1 The way to get the general cosmological solution of the CCM